OpenStudy (anonymous):

If f(x) = −3x − 3 and g(x) = 2x + 1, find f over g(−2).

OpenStudy (anonymous):

Choices: −3 3 −1 1

OpenStudy (anonymous):

\[f(x)=ax\] \[g(x)=bx\] \[\frac{f}{g}(x)=\frac{f(x)}{g(x)}\] \[\frac{f(x)}{g(x)}=\frac{ax}{bx}\]

OpenStudy (anonymous):

Use that information I gave you to solve your current question.